OpenStudy (b77w):

Why is there no x-intercept for this equation? How do I show it algebraically? y=log4(4x+8)+1

OpenStudy (anonymous):

\(x\) intercept mean where \(y=0\) set \[\log_4(4x+8)+1=0\] and solve for \(x\)

OpenStudy (b77w):

Yeah I get 4^-2.5

OpenStudy (anonymous):

frankly i don't ee what that cannot be done

OpenStudy (b77w):

The problem is that I keep getting an answer when the teacher said there isn't one

OpenStudy (ranga):

log(A) is not defined for A <= 0

OpenStudy (anonymous):

\[\log_4(4x+8)+1=0\\ \log_4(4x+8)=-1\\ 4x+8=4^{-1}=\frac{1}{4}\] why not? look like you can solve it to me

OpenStudy (anonymous):

i get \(x=-\frac{31}{16}\)

OpenStudy (anonymous):

which is not the answer you got, but it is an answer your math teacher i wrong, this crosses the \(x\) axis for sure

OpenStudy (b77w):

thank you

OpenStudy (anonymous):

yw

OpenStudy (loser66):

@ranga I am with you , however, in this case A >0

OpenStudy (anonymous):

?

OpenStudy (ranga):

I did not do the calculation. I was looking at one possible reason why this function may not have a root.

OpenStudy (anonymous):

you cannot take the log of a negative number which means to the domain of this function is \(x\geq -2\)

OpenStudy (loser66):

4(-31/16 )+8 >0

OpenStudy (anonymous):

but the range of a logarithmic function is \((-\infty, \infty)\)

OpenStudy (isaiah.feynman):

Its undefined because the curve approaches the y axis but never touches it.

OpenStudy (isaiah.feynman):

Domain is x>0

OpenStudy (ranga):

In this case the domain is > -2. (-2, infinity)

OpenStudy (isaiah.feynman):

That agrees with @ranga 's statement.